OpenStudy (anonymous):

The area of a triangle whose height is 1 more than 6 times its base is 13 square feet. If the base of the triangle is x feet, which equation models this situation? 1) 6x^2+x-12=0 2) 6x^2+x-13=0 3) 6x^2+x-26=0 4) 36x^2+6x-26=0

OpenStudy (displayerror):

The area of a triangle is given by \[\text{Area} = \frac{1}{2} \times \text{Base} \times \text{Height}\] We need to develop expressions for the \(\text{Base}\) and \(\text{Height}\) terms. We know what the base is--we're told that it can be represented by \(\text{Base} = x\) We're also told that the height is "1 more than 6 times its base". Let's break this down into math. "1 more than" suggests addition. Think about this: wouldn't you say that 3 apples is 1 more than 2 apples? "6 times its base" suggests multiplication. After all, that's the expression we typically use to express multiplication. You should be able to create an expression for the \(\text{Height}\) term now. Plug all of your information into the expression for \(\text{Area}\) and simplify to get an expression similar to the choices provided.